# SMaRTT-REPS: Sender-based Marked Rapidly-adapting Trimmed & Timed Transport with Recycled Entropies

### v1

@@ Abstract.
With the rapid growth of machine learning (ML) workloads in datacenters, existing congestion control (CC) algorithms fail to deliver the required performance at scale. ML traffic is bursty and bulk-synchronous and thus requires quick reaction and strong fairness. We show that existing CC algorithms that use delay as a main signal react too slowly and are not always fair. We design SMaRTT, a simple sender-based CC algorithm that combines delay, ECN, and optional packet trimming for fast and precise window adjustments. At the core of SMaRTT lies the novel QuickAdapt algorithm that accurately estimates the bandwidth at the receiver. We show how to combine SMaRTT with a new per-packet traffic load-balancing algorithm called REPS to effectively reroute packets around congested hotspots as well as flaky or failing links. Our evaluation shows that SMaRTT alone outperforms EQDS, Swift, BBR, and MPRDMA by up to 50% on modern datacenter networks.
@@ 1.1.1. Fairness
We show this effect in Fig. 1(a), where we report the Cumulative Distribution Function (CDF) of the Flow Completion Time (FCT) of 2MiB flows in a permutation scenario with 1,024 nodes, for a 8:1 oversubscribed network (more details on the setup in Sec. 4). We can notice how, with Swift (Kumar et al., 2020), a delay-based algorithm, the FCT of the slowest flow is 18% longer than the fastest flow. An even more pronounced effect can be observed with BBR (Cardwell et al., 2017) (which we observed needs up to seven round trip times (RTTs) to adjust its sending rate). MPRDMA (Lu et al., 2018), a per-packet ECN based mechanism suffers from its inherit unfairness which is more visible for small messages. As we discuss further in Sec. 1.1.2, vanilla EQDS (Olteanu et al., 2022), although more fair, does not manage fabric congestion as effectively as sender-based CC algorithms. On the other hand, the algorithm we propose (SMaRTT) quickly and precisely reacts to congestion by relying on different congestion signals (delay, ECN markings, and packet trimming), improving fairness and overall performance. Finally, we also show how using SMaRTT together with EQDS can improve the performance of the latter (more details in Sec. 5.1).
@@ 1.1.2. Fabric Congestion Management
Many CC algorithms have been optimized for managing congestion at the final hop caused by incast (Handley et al., 2017; Olteanu et al., 2022) due to its relevance for traditional TCP-based workloads (Chen et al., 2009). However, with the emergence of new workloads, those algorithms provide sub-optimal performance, since they might have limited visibility of congestion happening in the network fabric. This type of congestion is often the result of oversubscription (i.e., traffic exceeding the capacity of the network), common at upper tiers of datacenter networks (Arslan et al., 2023; Gibson et al., 2022), but can also occur in a non-oversubscribed setting as a result of ECMP collisions (Alizadeh et al., 2014; Ghorbani et al., 2017; Cao et al., 2013). This is shown in Fig. 1(b), where we report the CDF of the FCT of 32MiB flows in a permutation scenario with 1,024 nodes, for an 8:1 oversubscribed network. In this scenario, flows are big enough to allow Swift and MPRDMA to react effectively to congestion. However, FCT with EQDS is higher due to its inability to adapt to fabric congestion, hence underutilizing network bandwidth due to trimmed packets and re-transmissions. As we show in Sec. 5.1, this can be improved by complementing EQDS with SMaRTT.
@@ 1.2. Contributions
To fulfill these requirements, we introduce a new CC algorithm which we call SMaRTT (Sender-based Marked Rapidly-adapting Trimmed & Timed Transport), together with a novel load balancing scheme called REPS (Recycled Entropy Packet Spraying). SMaRTT is a sender-based CC algorithm running over lossy Ethernet. To be easily deployable, SMaRTT does not rely on any advanced in-band network telemetry (INT). It instead leverages a combination of ECN-marking and round-trip-time (RTT) measurements, two signals that have been commonly used in a mutually exclusive manner despite each of them possessing distinctive valuable benefits that complement one another (as we show in Sec. 3.1). Last, SMaRTT efficiently uses memory and compute resources, since it only needs to store a few bytes per flow and does not require any pull queue or specific control packets, as common in many receiver-driven algorithms. Specifically, SMaRTT-REPS makes the following key contributions:
@@
A novel mechanism that can quickly converge to the achievable bandwidth capacity at any bottleneck, typically within one RTT of receiving a severe congestion signal (Sec. 3.3). QuickAdapt relies on packet trimming or, in the case where it is not supported, on timeouts (with minimal performance impact, Sec. 4.2). Using QuickAdapt, SMaRTT can react to severe congestion as quickly as a receiver-based protocol but without requiring any special support on the receiver side (e.g., pull queues). We also argue that this feature can be retrofitted to any transport mechanism operating within tight RTT bounds, thus significantly reducing tail latencies. In particular, QuickAdapt can reduce queue occupancy much earlier than other sender-based mechanisms.
@@
A new congestion window management technique to improve bandwidth fairness (Sec. 3.2.1 and Sec. 3.2.3). In a nutshell, the congestion window is increased or decreased more conservatively (and proportionally to the currently acknowledged packet size) when the ECN and delay congestion signal disagree on the presence of congestion. To reduce any small degree of unfairness that might be present, SMaRTT rapidly increases the congestion window of flows (Sec. 3.4) if no congestion is detected for a period of time.
- REPS:
A powerful and lightweight adaptive load balancing mechanism called Recycled Entropy Packet Spraying (REPS, Sec. 3.6). REPS can sustain near-optimal performance at high bandwidth utilization levels and effectively route around path blackholes. Unlike complex in-network load balancing implementations (De Sensi et al., 2019; Rocher-González et al., 2022), REPS does not require any switch support besides ECMP. We demonstrate how SMaRTT-REPS effectively orchestrates congestion control and adaptive load balancing. On one side, the CC algorithm does not inhibit early yet mild congestion signals, which are crucial for load balancing. On the other side, load balancing does not aggressively shift substantial portions of traffic away to different paths, which could make the received congestion signals irrelevant. We demonstrate REPS in a real deployment on 128 nodes connected through a two-tier fat-tree network. ‘
@@ 2.1. Congestion Signals
SMaRTT uses both ECN and delay as congestion signals to react quickly and accurately to congestion. To efficiently manage severe congestion scenarios, SMaRTT relies on trimming or, if it is not supported, on timeouts. SMaRTT does not rely on complex in-network telemetry.
@@ 2.2. Rate Control
SMaRTT is a sender-based mechanism that, thanks to our novel QuickAdapt algorithm (Sec. 3.3), can precisely regulate the transmission rate in incast scenarios. At the same time, SMaRTT is resource-efficient since it does not require pull queues. As already discussed, to improve deployability, SMaRTT does not rely on any complex or new feature to be implemented in switches.
3. SMaRTT Design
@@
SMaRTT is a high-performance, easy-to-deploy, and lightweight CC algorithm that can quickly react to congestion happening both at the last hop (incast) and in any other part of the network. SMaRTT keeps a number of in-flight bytes equal to the size of a congestion window (cwnd). SMaRTT updates the window depending on the estimated network congestion, striving to keep packets RTT lower than a pre-determined and constant target RTT. This can be expressed relatively to the base RTT (i.e., the RTT on an empty network). The base RTT be computed when the datacenter network is set up (once for each possible hop count) or recomputed dynamically like what is done by Swift (Kumar et al., 2020). In the rest of the paper, we assume a target RTT equal to x the base RTT. Consequently, we can also easily compute a specific target RTT for each hop count by setting it to 50% more than the base RTT (details in Sec. 3.5).
@@
SMaRTT relies on widely available features like ECMP, ECN marking, and packet trimming. ECMP and ECN marking are available on most datacenter switches (Broadcom, 2024a; Nvidia, 2024; Hoefler et al., 2023), whereas packet trimming can be implemented on existing switches (Adrian et al., 2022). If ECMP is not supported, we can still run the CC algorithm without the REPS load balancing. If packet trimming is not supported, SMaRTT falls back to timeouts, with limited performance impact (Sec. 4.2). Similarly to EQDS (Olteanu et al., 2022), we assume that each switch port has a high-priority queue to forward control packets such as trims and ACKs, and a lower-priority queue to forward data packets. By doing so, packets carrying potential congestion information are prioritized so that SMaRTT can react quickly.
@@
This work mainly focuses on widely used fat tree networks (Poutievski et al., 2022; De Sensi et al., 2022). However, SMaRTT could be easily adapted to work with other topologies such as torus, Dragonfly (Kim et al., 2008; Gibson et al., 2022; De Sensi et al., 2020), BCube (Guo et al., 2009), SlimFly (Besta and Hoefler, 2014), HammingMesh (Hoefler et al., 2022) and others. Moreover, we intentionally limit buffering to a worst-case scenario, assuming that each per-port queue is one bandwidth-delay product (BDP) in size.
@@
In the following, we motivate the choice of using both delay and ECN as congestion signal (Sec. 3.1), and then describe the main control loop (Sec. 3.2). We then present QuickAdapt (Sec. 3.3) and FastIncrease (Sec. 3.4), two techniques we introduce to further improve SMaRTT reactiveness. Eventually, we discuss how to select SMaRTT parameters (Sec. 3.5), and how we couple it with the novel REPS load balancing algorithm (Sec. 3.6).
3.1. SMaRTT Congestion Signals
Choosing the congestion signal is critical for the whole design of the protocol, and SMaRTT uses both ECN and delay (RTT). SMaRTT can also leverage packet trimming to react to severe congestion, which we discuss in Sec. 3.3. In Fig. 2, we compare the full version of SMaRTT with a simple CC using ECN and delay, by simulating 8:1 incast traffic on 1,024 nodes fat tree with 800Gbps links. In all the simulation results, we use the htsim packet-level network simulator (Handley et al., 2017). We use the same congestion control logic for both ECN only and RTT only (we decrease the congestion window by half an MTU per packet at most in response to high delay or ECN marking). We observe that the congestion windows decrease faster when using ECN since, when using delay, the receiver gets the congestion signal only after the packet crosses the entire queue. On the other hand, RTT only is more fair. Finally we show SMaRTT as reference. With it, we take advantage of both signals and also use QuickAdapt (Sec /3.3) to immediately adjust its windows to the right value.
@@ 3.1. SMaRTT Congestion Signals
Last, it is important to determine how often to react to congestion signals. Many algorithms like MPRDMA react once per ACK to improve responsiveness. However, it might be hard to keep up when increasing network bandwidth. For this reason, other algorithms like Swift or DCQCN react once per period of time (e.g., once per RTT). As shown in Fig. 3(b), reacting on a per-packet basis can result in slightly better performance. Indeed, the congestion window size is adapted more smoothly, improving reaction times and reducing unfairness. We also observe how, even reacting only once every 50 packets does not significantly impact SMaRTT performance, which is within the 5% of the per-packet reaction scenario.
@@ 3.2. Main Control Loop
In SMaRTT, the sender recomputes the congestion window size every time it receives an ACK. When an ACK is received, four different scenarios can occur depending on the severity of congestion (or lack thereof), as reported by the two congestion signals: Fair Decrease (Sec. 3.2.1), Multiplicative Decrease (Sec. 3.2.2), Fair Increase (Sec. 3.2.3), Multiplicative Increase (Sec. 3.2.4).
@@ 3.2.1. Fair Decrease (FD)
The ACK is ECN-marked, but the reported RTT is smaller than the target one. In this case, queues are building up, but congestion has not yet affected the acknowledged packet. SMaRTT proactively reacts by gently decreasing the window regardless of the packet’s RTT. To improve fairness (see Sec. 4.3), the congestion window is decreased more aggressively for flows with larger windows (and thus, likely, more in-flight bytes) as follows:
@@ 3.2.2. Multiplicative Decrease (MD)
The ACK is ECN-marked and reports an RTT higher than the target. This indicates a worst-case scenario and, as a consequence, SMaRTT rapidly reduces the congestion window proportionally to the difference between the target and measured RTT, as follows:
where and are the target and measured RTT, respectively. SMaRTT additionally applies Fair Decrease to improve fairness (Sec. 3.2.1, not shown explicitly in Eq. 2). In any case, the multiplicative decrease reduces at most by the acknowledged packet’s size. is a multiplicative decrease constant selected so that when the RTT is twice the target RTT, the congestion window is decreased aggressively by a packet for each ACK. Thus, we want to have , which implies .
@@ 3.2.3. Fair Increase (FI)
The ACK is not ECN-marked but reports an RTT higher than the target one. This might indicate that the delay signal is outdated and the queues are not growing significantly. Consequently, SMaRTT gently increases the congestion window, ignoring the delay signal. To improve fairness (see Sec. 4.3), the window is increased more aggressively for flows with smaller windows as follows:
@@ 3.2.4. Multiplicative Increase (MI)
The received ACK is not ECN-marked and reports an RTT smaller than the target one. In this case, SMaRTT increases the congestion window proportionally to the difference between the target and measured RTTs, as follows:
is a multiplicative increase constant selected to increase the congestion window by at most one MTU per RTT. If we assume ACKs are received in an RTT, SMaRTT would increase the window size by per RTT. Because we want this quantity to be smaller than an MTU, . Since , where is the base RTT, we set . This multiplicative increase is followed by a fair increase to improve fairness (Sec. 3.2.3, not shown explicitly in Eq. 4).

### v3

# FASTFLOW: Flexible Adaptive Congestion Control for High-Performance Datacenters

@@ Abstract.
The increasing demand of machine learning (ML) workloads in datacenters places significant stress on current congestion control (CC) algorithms, many of which struggle to maintain performance at scale. These workloads generate bursty, synchronized traffic that requires both rapid response and fairness across flows. Unfortunately, existing CC algorithms that rely heavily on delay as a primary congestion signal often fail to react quickly enough and do not consistently ensure fairness. In this paper, we propose FASTFLOW, a streamlined sender-based CC algorithm that integrates delay, ECN signals, and optional packet trimming to achieve precise, real-time adjustments to congestion windows. Central to FASTFLOW is the QuickAdapt mechanism, which provides accurate bandwidth estimation at the receiver, enabling faster reactions to network conditions. We also show that FASTFLOW can effectively enhance receiver-based algorithms such as EQDS by improving their ability to manage in-network congestion. Our evaluation reveals that FASTFLOW outperforms cutting-edge solutions, including EQDS, Swift, BBR, and MPRDMA, delivering up to 50% performance improvements in modern datacenter networks.
@@ 1.1.1. Fairness
We show this effect in Fig. 1(a), where we report the Cumulative Distribution Function (CDF) of the Flow Completion Time (FCT) of 2MiB flows in a permutation scenario with 1,024 nodes, for a 8:1 oversubscribed network (more details on the setup in Sec. 4). We can notice how, with Swift (Kumar et al., 2020), a delay-based algorithm, the FCT of the slowest flow is 18% longer than the fastest flow. An even more pronounced effect can be observed with BBR (Cardwell et al., 2017) (which we observed needs up to seven round trip times (RTTs) to adjust its sending rate). MPRDMA (Lu et al., 2018), a per-packet ECN based mechanism suffers from its inherit unfairness which is more visible for small messages. Last, as we discuss further in Sec. 1.1.2, vanilla EQDS (Olteanu et al., 2022), although more fair, does not manage fabric congestion as effectively as sender-based CC algorithms. On the other hand, the algorithm we propose (FASTFLOW) quickly and precisely reacts to congestion by relying on different congestion signals (delay, ECN markings, and packet trimming), improving fairness and overall performance.
@@ 1.1.2. Fabric Congestion Management
Many CC algorithms have been optimized for managing congestion at the final hop caused by incast (Handley et al., 2017; Olteanu et al., 2022) due to its relevance for traditional TCP-based workloads (Chen et al., 2009). However, with the emergence of new workloads, those algorithms provide sub-optimal performance, since they might have limited visibility of congestion happening in the network fabric. This type of congestion is often the result of oversubscription (i.e., traffic exceeding the capacity of the network), common at upper tiers of datacenter networks (Arslan et al., 2023; Gibson et al., 2022), but can also occur in a non-oversubscribed setting as a result of ECMP collisions (Alizadeh et al., 2014; Ghorbani et al., 2017; Cao et al., 2013). This is shown in Fig. 1(b), where we report the CDF of the FCT of 32MiB flows in a permutation scenario with 1,024 nodes, for an 8:1 oversubscribed network. In this scenario, flows are big enough to allow Swift and MPRDMA to react effectively to congestion. However, FCT with EQDS is higher due to its inability to adapt to fabric congestion, hence underutilizing network bandwidth due to trimmed packets and re-transmissions. As we show in Sec. 5.1, this can be improved by complementing EQDS with FASTFLOW.
@@ 1.2. Contributions
To fulfill these requirements, we introduce a new CC algorithm which we call FASTFLOW (Sender-based Marked Rapidly-adapting Trimmed & Timed Transport). FASTFLOW is a sender-based CC algorithm running over lossy Ethernet. To be easily deployable, FASTFLOW does not rely on any advanced in-band network telemetry (INT). It instead leverages a combination of ECN-marking and round-trip-time (RTT) measurements, two signals that have been commonly used in a mutually exclusive manner despite each of them possessing distinctive valuable benefits that complement one another (as we show in Sec. 3.1). Last, FASTFLOW efficiently uses memory and compute resources, since it only needs to store a few bytes per flow and does not require any pull queue or specific control packets, as common in many receiver-driven algorithms. Specifically, FASTFLOW makes the following key contributions:
@@
A novel mechanism that can quickly converge to the achievable bandwidth capacity at any bottleneck, typically within one RTT of receiving a severe congestion signal (Sec. 3.3). QuickAdapt relies on packet trimming or, in the case where it is not supported, on timeouts (with minimal performance impact, Sec. 4.1). Using QuickAdapt, FASTFLOW can react to severe congestion as quickly as a receiver-based protocol but without requiring any special support on the receiver side (e.g., pull queues). We also argue that this feature can be retrofitted to any transport mechanism operating within tight RTT bounds, thus significantly reducing tail latencies. In particular, QuickAdapt can reduce queue occupancy much earlier than other sender-based mechanisms.
@@
A new congestion window management technique to improve bandwidth fairness (Sec. 3.2.1 and Sec. 3.2.3). In a nutshell, the congestion window is increased or decreased more conservatively (and proportionally to the currently acknowledged packet size) when the ECN and delay congestion signal disagree on the presence of congestion. To reduce any small degree of unfairness that might be present, FASTFLOW rapidly increases the congestion window of flows (Sec. 3.4) if no congestion is detected for a period of time.
@@ 2.1. Congestion Signals
FASTFLOW uses both ECN and delay as congestion signals to react quickly and accurately to congestion. To efficiently manage severe congestion scenarios, FASTFLOW relies on trimming or, if it is not supported, on timeouts. To enable FASTFLOW to be deployed easily on existing networks, FASTFLOW does not rely on complex in-network telemetry.
@@ 2.2. Rate Control
FASTFLOW is a sender-based mechanism that, thanks to our novel QuickAdapt algorithm (Sec. 3.3), can precisely regulate the transmission rate in incast scenarios. At the same time, FASTFLOW is resource-efficient since it does not require pull queues. As already discussed, to improve deployability, FASTFLOW does not rely on any complex or new feature to be implemented in switches.
3. FASTFLOW Design
@@
FASTFLOW is a high-performance, easy-to-deploy, and lightweight CC algorithm that can quickly react to congestion happening both at the last hop (incast) and in any other part of the network. FASTFLOW keeps a number of in-flight bytes equal to the size of a congestion window (cwnd). FASTFLOW updates the window depending on the estimated network congestion, striving to keep packets RTT lower than a pre-determined and constant target RTT. This can be expressed relatively to the base RTT (i.e., the RTT on an empty network). The base RTT be computed when the datacenter network is set up (once for each possible hop count) or recomputed dynamically like what is done by Swift (Kumar et al., 2020). In the rest of the paper, we assume a target RTT equal to x the base RTT. Consequently, we can also easily compute a specific target RTT for each hop count by setting it to 50% more than the base RTT (details in Sec. 3.5).
@@
FASTFLOW relies on widely available features like ECMP, ECN marking, and packet trimming. ECMP and ECN marking are available on most datacenter switches (Broadcom, 2024a; Nvidia, 2024; Hoefler et al., 2023), whereas packet trimming can be implemented on existing switches (Adrian et al., 2022). If ECMP is not supported, we can still run the CC algorithm. If packet trimming is not supported, FASTFLOW falls back to timeouts, with limited performance impact (Sec. 4.1). Similarly to EQDS (Olteanu et al., 2022), we assume that each switch port has a high-priority queue to forward control packets such as trims and ACKs, and a lower-priority queue to forward data packets. By doing so, packets carrying potential congestion information are prioritized so that FASTFLOW can react quickly.
@@
This work mainly focuses on widely used fat tree networks (Poutievski et al., 2022; De Sensi et al., 2022). However, FASTFLOW could be easily adapted to work with other topologies such as torus, Dragonfly (Kim et al., 2008; Gibson et al., 2022; De Sensi et al., 2020), BCube (Guo et al., 2009), SlimFly (Besta and Hoefler, 2014), HammingMesh (Hoefler et al., 2022) and others. Moreover, we intentionally limit buffering to a worst-case scenario, assuming that each per-port queue is one bandwidth-delay product (BDP) in size.
@@
In the following, we motivate the choice of using both delay and ECN as congestion signal (Sec. 3.1), and then describe the main control loop (Sec. 3.2). We then present QuickAdapt (Sec. 3.3) and FastIncrease (Sec. 3.4), two techniques we introduce to further improve FASTFLOW reactiveness. Eventually, we discuss how to select FASTFLOW parameters (Sec. 3.5).
3.1. FASTFLOW Congestion Signals
Choosing the congestion signal is critical for the whole design of the protocol, and FASTFLOW uses both ECN and delay (RTT). FASTFLOW can also leverage packet trimming to react to severe congestion, which we discuss in Sec. 3.3. In Fig. 2, we compare the full version of FASTFLOW with a simple CC using ECN and delay, by simulating 8:1 incast traffic on 1,024 nodes fat tree with 800Gbps links. In all the simulation results, we use the htsim packet-level network simulator (Handley et al., 2017). We use the same congestion control logic for both ECN only and RTT only (we decrease the congestion window by half an MTU per packet at most in response to high delay or ECN marking). We observe that the congestion windows decrease faster when using ECN since, when using delay, the receiver gets the congestion signal only after the packet crosses the entire queue. On the other hand, RTT only is more fair. Finally we show FASTFLOW as reference. With it, we take advantage of both signals and also use QuickAdapt (Sec /3.3) to immediately adjust its windows to the right value.
@@ 3.1. FASTFLOW Congestion Signals
Last, it is important to determine how often to react to congestion signals. Many algorithms like MPRDMA react once per ACK to improve responsiveness. However, it might be hard to keep up when increasing network bandwidth. For this reason, other algorithms like Swift or DCQCN react once per period of time (e.g., once per RTT). As shown in Fig. 3(b), reacting on a per-packet basis can result in slightly better performance. Indeed, the congestion window size is adapted more smoothly, improving reaction times and reducing unfairness. We also observe how, even reacting only once every 50 packets does not significantly impact FASTFLOW performance, which is within the 5% of the per-packet reaction scenario.
@@ 3.2. Main Control Loop
In FASTFLOW, the sender recomputes the congestion window size every time it receives an ACK. When an ACK is received, four different scenarios can occur depending on the severity of congestion (or lack thereof), as reported by the two congestion signals: Fair Decrease (Sec. 3.2.1), Multiplicative Decrease (Sec. 3.2.2), Fair Increase (Sec. 3.2.3), Multiplicative Increase (Sec. 3.2.4).
@@ 3.2.1. Fair Decrease (FD)
The ACK is ECN-marked, but the reported RTT is smaller than the target one. In this case, queues are building up, but congestion has not yet affected the acknowledged packet. FASTFLOW proactively reacts by gently decreasing the window regardless of the packet’s RTT. To improve fairness (see Sec. 4.2), the congestion window is decreased more aggressively for flows with larger windows (and thus, likely, more in-flight bytes) as follows:
@@ 3.2.2. Multiplicative Decrease (MD)
The ACK is ECN-marked and reports an RTT higher than the target. This indicates a worst-case scenario and, as a consequence, FASTFLOW rapidly reduces the congestion window proportionally to the difference between the target and measured RTT, as follows:
where and are the target and measured RTT, respectively. FASTFLOW additionally applies Fair Decrease to improve fairness (Sec. 3.2.1, not shown explicitly in Eq. 2). In any case, the multiplicative decrease reduces at most by the acknowledged packet’s size. is a multiplicative decrease constant selected so that when the RTT is twice the target RTT, the congestion window is decreased aggressively by a packet for each ACK. Thus, we want to have , which implies .
@@ 3.2.3. Fair Increase (FI)
The ACK is not ECN-marked but reports an RTT higher than the target one. This might indicate that the delay signal is outdated and the queues are not growing significantly. Consequently, FASTFLOW gently increases the congestion window, ignoring the delay signal. To improve fairness (see Sec. 4.2), the window is increased more aggressively for flows with smaller windows as follows:
@@ 3.2.4. Multiplicative Increase (MI)
The received ACK is not ECN-marked and reports an RTT smaller than the target one. In this case, FASTFLOW increases the congestion window proportionally to the difference between the target and measured RTTs, as follows:
is a multiplicative increase constant selected to increase the congestion window by at most one MTU per RTT. If we assume ACKs are received in an RTT, FASTFLOW would increase the window size by per RTT. Because we want this quantity to be smaller than an MTU, . Since , where is the base RTT, we set . This multiplicative increase is followed by a fair increase to improve fairness (Sec. 3.2.3, not shown explicitly in Eq. 4).